A train starts at rest travels 30m N after 20m. The trains stop after traveling 10m more S in 5min

a. What is the d traveled by train?

b. What is the d of the train?

c. What is the speed of the train in m/s?

d. What is the v of the train in m/s?




1
Expert's answer
2021-09-28T11:11:30-0400


a) the distance


"l=l_1+l_2=\rm 30\: m+10\: m=40\: m"

b) the displacement


"d=l_1-l_2=\rm 30\: m-10\: m=20\: m \quad North"

c) the speed

"v=\frac{l}{t}=\rm \frac{40\: m}{25\: min}=\frac{40\: m}{25\: min}=1.6 \: m/min"

d) the velocity

"v=\frac{d}{t}=\rm \frac{20\: m}{25\: min}=0.8\: m/min \quad North"